OpenStudy (anonymous):

what is the equation in standard form of the line which passes through (4,-2) and has a slope of -3?

OpenStudy (anonymous):

A.) x+3y=10 B.) x+3y=-10 C.) 3x+y=10 D.) 3x+y=-10

OpenStudy (midhun.madhu1987):

The equation is given by y - y1 = m ( x- x1) m= slope just plug in the values...

OpenStudy (anonymous):

y= mx+ c, slope of -3 is m(gradient) so y= -3x+c to find c(x- intercept), sub in the values (4, -2) -2= -3(4)+ c so c= 10 therefore the answer is y= -3x+ 10